The cost (in dollars) of producing x units of a certain commodity is C(x)=500+10x+0.05x^2

I) find the average rate of change of C with respect x when the production level is from x=100 to x=105 units
ii) find the instantaneous rate of change of C with respect to x=100
1
Expert's answer
2020-05-20T18:59:24-0400

Given "C(x)=500+10x+0.05x^2"

i) The average rate of change is the change in "y" value over the change in "x" value for two distinct points, hence if "x" changes from 100 to 105, then you may evaluate the average rate of change of "C(x)"  such that: "\frac{C(105)-C(100)}{105 - 100 }" .

Now, "C(105)-C(100) = (500 + 10(105) + 0.05 (105)^2) - (500 + 10(100) + 0.05 (100)^2) = (1050 - 1000) + (551.25 - 500) = 101.25"

So, Average rate of change of "C(x)" from 100 to 105 is "\frac{101.25}{5} = 20.25".


ii)  The instantaneous rate of change of "C" with respect to x=100 is "C'(x)|_{x=100}" .

Now, given "C(x)=500+10x+0.05x^2"

"\implies C'(x) = 10 + 0.1 x"

The instantaneous rate of change of "C" with respect to 100 is "C'(100) = 10 + 0.1(100) = 20" .